# Democratizing Uncertainty Quantification

### v2

@@ B.4.1 L2-Sea optimization
where the objective function $f$ is the total resistance $R$, the equality constraints $h_{i}$ (automatically satisfied by the shape modification method) include ($h_{1}$) fixed length between perpendiculars of the vessel and ($h_{2}$) fixed vessel displacement, whereas the inequality constraints $g_{1},\ldots,g_{4}$ encode geometrical constrains on the modified hull in relation to the original one. In particular, $g_{1}$ relates to the beam, $g_{2}$ to the draft (both allowing ±$\pm 5\%$ of maximum variation), and $g_{3},g_{4}$ relate to the minimum cylindrical dimensions to contain the sonar in the bow dome, see [123] for details: these are the four additional outputs of the L2-sea model that were hinted at the end of Section B.1.2.

### v4

@@ A.4.1 L2-Sea optimization
where the objective function $f$ is the total resistance $R$, the equality constraints $h_{i}$ (automatically satisfied by the shape modification method) include ($h_{1}$) fixed length between perpendiculars of the vessel and ($h_{2}$) fixed vessel displacement, whereas the inequality constraints $g_{1},\ldots,g_{4}$ encode geometrical constrains on the modified hull in relation to the original one. In particular, $g_{1}$ relates to the beam, $g_{2}$ to the draft (both allowing ±$\pm 5\%$ of maximum variation), and $g_{3},g_{4}$ relate to the minimum cylindrical dimensions to contain the sonar in the bow dome, see [67] for details: these are the four additional outputs of the L2-sea model that were hinted at the end of Section A.1.2.